### v1

# Full-Duplex-Bench: A Benchmark to Evaluate Full-duplex Spoken Dialogue Models on Turn-taking Capabilities

Abstract
Spoken dialogue modeling introduces unique challenges beyond text-based language modeling, demanding robust turn-taking, backchanneling, and real-time interaction. Although most Spoken Dialogue Models (SDMs) rely on half-duplex processing (handling speech one turn at a time), emerging full-duplex SDMs can listen and speak simultaneously, enabling more natural and engaging conversations. However, current evaluations of such models remain limited, often focusing on turn-based metrics or high-level corpus analyses (e.g., turn gaps, pauses). To address this gap, we present Full-Duplex-Bench [^1] Data, codebase, and demo samples will be released here https://github.com/DanielLin94144/Full-Duplex-Bench., a new benchmark that systematically evaluates key conversational behaviors: pause handling, backchanneling, turn-taking, and interruption management. Our framework uses automatic metrics for consistent and reproducible assessments of SDMs’ interactive performance. By offering an open and standardized evaluation benchmark, we aim to advance spoken dialogue modeling and encourage the development of more interactive and natural dialogue systems.
Full-Duplex-Bench: A Benchmark to Evaluate Full-duplex Spoken Dialogue Models on Turn-taking Capabilities
Guan-Ting Lin1  Jiachen Lian2thanks: Equal contribution, listed alphabetically.  Tingle Li2∗ Qirui Wang3∗ Gopala Anumanchipalli2  Alexander H. Liu4  Hung-yi Lee1 1Graduate Institute of Communication Engineering, National Taiwan University 2UC Berkeley  3University of Washington  4MIT CSAIL
1 Introduction
Natural spoken dialogue is characterized by complex dynamics (Schegloff, 2000; Gravano and Hirschberg, 2011). Different from text-based language modeling, spoken dialogue involves problems beyond language understanding for communication, including turn-taking (Duncan, 1972; Gravano and Hirschberg, 2011; Raux and Eskenazi, 2012), and backchanneling (Schegloff, 1982). These fundamental aspects facilitate mutual understanding, engagement, and social connection. Over the years, Spoken Dialogue Models (SDMs) have become a major research focus, particularly with the rise of voice assistants such as Apple Siri, Google Assistant, and Amazon Alexa.
Figure 1: Overview pipeline of Full-Duplex-Bench. We feed user audio streams to an open-source full-duplex SDM, which produces time-synchronous output. We then perform post-processing to align both streams at the transcript level, enabling automatic evaluation along multiple dimensions.
Existing SDMs can be broadly categorized into two types: (1) Half-duplex SDMs: These models operate based on a turn-by-turn protocol, processing one audio stream at a time and switching roles in response to turn-changing signals. (2) Full-duplex SDMs: These models can simultaneously listen and speak, allowing them to capture the subtle timing of spoken conversation. They process continuous audio streams and model overlapping speech, pauses, and background noise. Furthermore, these models can capture essential spoken interaction behaviors, such as backchanneling, offering more natural and context-aware responses.
Although half-duplex SDMs have been the dominant approach, they often sound less natural than human speakers, who seamlessly listen and speak at the same time. Recent progress in models like GPT-4o voice mode has sparked an interest in full-duplex capabilities for more human-like dialogues. Consequently, many new SDMs have emerged (Veluri et al., 2024; Défossez et al., 2024; Wang et al., 2024a; Fu et al., 2024; Wang et al., 2024c; Chen et al., 2025; Ji et al., 2024), with different architectures to enable real-time communication. As an increasing number of these systems are proposed, fair and open evaluation is critical to guide future advances in spoken dialogue research.
Evaluating SDMs is challenging because human conversation requires understanding a wide range of information and interpreting complex interactions within speech signals. Previous benchmarks have largely focused on: Content-based evaluation, measuring performance on tasks such as spoken question answering (Hassid et al., 2024; Nachmani et al., 2024; Shih et al., 2024; Lin et al., 2024c, 2022); Instruction-following evaluation, assessing a system’s ability to follow directives (Huang et al., 2024b, a); and Paralinguistic-based evaluation, examining factors like emotion or speaking style (Cheng et al., 2025; Ao et al., 2024; Yang et al., 2024; Lin et al., 2024a, b; Lin and Lee, 2024). However, these benchmarks mostly assume turn-based interactions, leaving the real-time aspects of full-duplex models under-explored.
Some recent studies attempt to evaluate interaction timing in full-duplex dialogue. dGSLM (Nguyen et al., 2023) examines voice activity patterns by comparing model outputs to ground truth gaps, interpausal units, and pauses. However, the corpus-level statistics are difficult to interpret, and the ground truth is based on a specific dialogue dataset. Recently, Talking-Turns (Arora et al., 2025) uses a specialized judge model to evaluate alignment with conversational behaviors, but the model is trained on a specific dialogue dataset, which may not generalize to different dialogues. While these works offer valuable insights, they emphasize the need for a more unified and reproducible approach to testing full-duplex SDMs in real-life scenarios.
In this paper, we address this gap by introducing the first benchmark designed to systematically evaluate key turn-taking behaviors in full-duplex SDMs (Figure 1). Our framework focuses on four crucial aspects – pause handling, backchanneling, turn-taking, and user interruption management, and uses automatic metrics for reliable comparisons. By evaluating our benchmark to SDMs, we can better understand how these systems handle real-time interactive dynamics. We will publicly release the benchmark and the associated codebase to encourage the further development of more engaging and conversational SDMs.
2 Related Works
2.1 Full-duplex Spoken Dialogue Models
Full-duplex SDMs aim to mimic the natural flow of human conversation by allowing simultaneous listening and speaking. Recent approaches can be divided into two main categories:
End-to-End Two-Channel Models
dGSLM (Nguyen et al., 2023) applies a Siamese network with cross-attention to handle two-channel spoken dialogue. Moshi (Défossez et al., 2024) uses parallel processing streams for speaker and listener speech. SyncLLM (Veluri et al., 2024) introduces a time-synchronous mechanism for autoregressive decoding across both channels. OmniFlatten (Zhang et al., 2024a) employs progressive post-training to jointly process flattened sequences of speech and text tokens, while SALMONN-omni (Yu et al., 2024) and MinMo (Chen et al., 2025) augment large language models with additional state tokens to refine speech-based intelligence. Parrot (Wang et al., 2025) uses next-token-pair prediction for two-channel modeling. RTTL-DG (Mai and Carson-Berndsen, 2025) leverages two-channel input with a dialogue manager to predict the speaking or listening state to determine when to generate the response.
Cascaded Models
Wang et al. (2024b) introduce control, speak, and listen tokens on top of a large language model with perception and motor modules, although their design still relies on speech-to-text and text-to-speech components. MiniCPM-Duplex (Zhang et al., 2024b) and MiniCPM-Duo (Xu et al., 2024) incorporate time-sliced chunks of text tokens, adapting time-synchronous methods for language modeling. VITA (Fu et al., 2024) and Freeze-Omni (Wang et al., 2024c) accept raw speech input directly, improving inference speed but still requiring text in the loop. These varied strategies highlight a clear need for standardized and open benchmarks that can evaluate turn-taking and real-time interaction across emerging full-duplex SDMs.
2.2 Evaluation Benchmarks
Benchmarks offer a common framework for comparing models and understanding their performance. In the speech community, several benchmarks have been proposed to evaluate specific aspects of speech models. The SUPERB series (wen Yang et al., 2021; Lin et al., 2023; Tsai et al., 2022) evaluates self-supervised speech representations across a variety of tasks, while ML-SUPERB (Shi et al., 2023), IndicSUPERB (Javed et al., 2023), and LeBenchmark (Evain et al., 2021) extend these evaluations to multilingual and non-English models.
Recent advances in universal spoken language models have resulted in benchmarks that evaluate both understanding and generation. These can be grouped as follows: Multi-task Instruction Following: Dynamic-SUPERB (Huang et al., 2024b, a), Voicebench (Chen et al., 2024), and AIR-Bench (Yang et al., 2024) test a model’s ability to perform a range of tasks based on specific instructions. Semantic Understanding: For low-level language skills such as syntax and grammar, benchmarks like Zerospeech (Nguyen et al., 2020) and datasets such as Spoken-StoryCloze (Mostafazadeh et al., 2017) have been introduced. Higher-level comprehension, including spoken question answering, is measured by benchmarks like Llama-Question (Nachmani et al., 2024) and Spoken WebQuestion. Paralinguistic Perception: VoxDialogue (Cheng et al., 2025), SD-Eval (Ao et al., 2024), and StyleTalk (Lin et al., 2024a) evaluate the ability to produce natural responses by considering features such as emotion, gender, and accent.
Despite these developments, no existing benchmark explicitly evaluates the turn-taking abilities of SDMs, a critical aspect of natural spoken interaction. To address this gap, we introduce a new benchmark that measures SDMs’ turn-taking behavior across multiple dimensions. This work propose a new benchmark for evaluating conversational fluidity, paving the way for the development of more interactive and natural dialogue systems.
3 Full-Duplex-Bench Framework
3.1 Overview
Our methodology utilizes unified speech input to simulate real-time user interactions with SDMs, creating controlled scenarios for each evaluation dimension. We prepare specific test samples for each aspect, collect the model’s speech responses, and apply CrispyWhisper (Zusag et al., 2024) to generate time-aligned transcriptions. Specialized metrics are then employed to assess performance in each dimension.
3.2 Prerequisite
For clarity, we first define several key terms that will be used in this paper:
Figure 2: Illustration of the four evaluation dimensions in Full-Duplex-Bench. (1) Pause Handling: the model stays silent during user pauses; (2) Backchanneling: the model offers short, timely acknowledgments; (3) Smooth Turn-taking: the model takes the turn in time; and (4) User Interruption: the model handles sudden user input with appropriate, well-timed responses.
Backchannel
In natural conversations, listeners often provide short verbal or nonverbal signals, known as backchannels, to indicate active engagement and understanding. These backchannels include utterances such as “mm-hmm,” and “uh-huh", which typically occur while another person is speaking. Effective backchanneling enhances conversational fluidity by signaling attentiveness without disrupting the speaker. Generally, backchanneling refers to short utterances produced by the listener while the speaker is talking.
In this work, we classify a speech segment as backchanneling if it meets the following criteria: (1) it has a short duration of less than 1 second; and (2) it contains fewer than two words. This ensures that speech is delivered at a reasonable pace and that brief utterances do not interrupt the current speaker’s turn. Although the concept of backchanneling differs across the literature (Ward and Tsukahara, 2000; Reece et al., 2023), developing a comprehensive detector remains a goal for future work.
Takeover (TO)
A takeover occurs when the model effectively assumes control of the conversation, dominating the turn and granting minimal opportunity for the user to speak. In this work, takeover is treated as a binary variable. If the model merely responds with silence or a backchannel, no takeover is deemed to have occurred. Conversely, any other non-silent speech that is not a backchannel indicates the model’s attempt to take over. Formally,
$\text{TO}=\begin{cases}0,&\text{if silence or backchannel}\\ 1,&\text{otherwise}\end{cases}$
Takeover Rate (TOR)
To quantify how often takeovers occur, we define the TOR as the average value of the binary TO variable across the dataset: $\text{TOR}=\frac{1}{N}\sum_{i=1}^{N}\text{TO}_{i}$, where $N$ is the total number of samples, and $\text{TO}_{i}$ is the binary takeover variable for sample $i$.
3.3 Evaluation Dimensions
Figure 2 illustrates the following four evaluation dimensions.
3.3.1 Pause Handling
Humans naturally pause and hesitate during conversations. Pauses can occur between consecutive turns or within the same sentence. These pauses are often not intended to yield the turn but rather to maintain control of the conversation. Therefore, taking over the turn during such pauses is undesirable and may lead to user dissatisfaction. Effective pause handling ensures that the model does not interrupt when a speaker takes a natural pause.
Research Question
Can the model recognize when the other speaker is still holding the turn and understand that it should not take over?
Metric
The ideal model behavior is to avoid taking over the turn while the user is speaking. To evaluate this, we use the Takeover Rate (TOR). A lower TOR signifies better pause management, indicating that the model effectively waits for the user’s turn to end. In contrast, a higher TOR suggests that the model is more likely to take over the conversation before the user has yielded the turn.
3.3.2 Backchanneling
Drawing on the definition in Section 3.2, we assess whether the model, when interacting with a dominant speaker, actively listens and provides backchannels at suitable moments to facilitate dialogue engagement. A model exhibiting human-like backchanneling behavior should respond at the right times and with an appropriate frequency.
Research Question
Can the model determine when to offer backchannels in a human-like manner without interrupting the speaker?
Metric
To measure how well models predict the timing of human backchannel cues (e.g., “uh-huh,” “mm-hmm”) during conversations, we use three metrics:
  • TOR: As with pause handling, the model should avoid dominating the turn, so a lower TOR is preferable.
  • Backchannel Frequency (Freq): Each backchannel event is counted and normalized by duration (events per second). When the model does not take over the turn (TOR = 0), a higher frequency suggests more effective backchanneling.
  • Jensen-Shannon Divergence (JSD): This captures the difference between the model’s predicted timing of backchannels and actual human timing. The model’s predictions form a probability distribution $P$, indicating the likelihood of a backchannel within each discrete time window. Similarly, the human ground truth forms a distribution $Q$ over the same set of time windows. We compute $\text{JSD}(P||Q)=\frac{1}{2}D_{\text{KL}}(P||M)+\frac{1}{2}D_{\text{KL}}(Q||M)$, where $M=\frac{P+Q}{2}$ is the average distribution, and $D_{\text{KL}}$ is the Kullback–Leibler divergence. Expanded, $\text{JSD}(P||Q)=\frac{1}{2}\sum_{i}P(i)\log\frac{P(i)}{M(i)}+\frac{1}{2}\sum_{i}Q(i)\log\frac{Q(i)}{M(i)}$. JSD ranges from 0 (perfect agreement) to 1 (complete divergence). A lower score indicates closer alignment with human backchannel timing. We only calculate this metric when the model does not take over the turn, where each backchannel event is counted as one-hot and normalized into a probability distribution. If the model stays silent throughout, we assume a uniform probability distribution, treating it as a random baseline without backchannel knowledge.
3.3.3 Smooth Turn Taking
Effective turn-taking is crucial for maintaining a natural and engaging conversation. In human dialogue, smooth turn transitions occur when speakers respond promptly without excessive delay or overlap. A well-designed model should be capable of recognizing turn boundaries and responding with appropriate timing to ensure fluid interactions.
Research Question
Can the model detect the end of a speaker’s turn and respond promptly without long pauses?
Metric
We measure the averaged response latency, the time (in seconds) between the end of the user’s speech and the start of the model’s response. Lower latency values indicate smoother turn-taking. In cases where the model fails to respond, we record the TOR. The latency is calculated only when TO equals 1. This avoids averaging with non-takeover periods, which would introduce significant variance due to periods of silence.
3.3.4 User Interruption
In human conversations, interruptions are common and can occur when a listener interjects mid-turn to clarify, disagree, or shift the discussion. A well-designed conversational model should be able to recognize and adapt to such interruptions by adjusting its response appropriately. Effective handling of interruptions ensures that the model remains responsive and coherent, even when the dialogue flow is disrupted.
Research Question
Can the model detect and adapt to user interruptions while maintaining a coherent and timely response?
Metric
The evaluation consists of three measures:
  • TOR: TOR is measured to ensure the model takes the turn following an interruption (ideally TOR = 1). The below metrics are calculated only when the TO is 1.
  • GPT-4o Score: A large language model evaluates the quality of the system’s response (coherence, relevance, and adaptability) on a scale from 0 to 5. Higher scores indicate better performance. Detailed instructions are provided in Appendix Table 5.
  • Latency After Interruption: The averaged time taken for the model to respond after an interruption, with lower averaged latency indicating smoother interaction.
Model Date E2E #ch Interrupt BC S2S Release
Transparent Models
dGSLM (Nguyen et al., 2023) 2022/3 2
FSM (Wang et al., 2024b) 2024/5 1 - -
MiniCPM-Duplex (Zhang et al., 2024b) 2024/6 1 - -
VITA (Fu et al., 2024) 2024/8 1 -
SyncLLM Veluri et al. (2024) 2024/9 2
Parrot Wang et al. (2025) 2024/9 2 -
MiniCPM-Duo (Xu et al., 2024) 2024/9 1 - -
Moshi (Défossez et al., 2024) 2024/10 2
SALMONN-omni (Yu et al., 2024) 2024/11 1 -
MinMo (Chen et al., 2025) 2025/1 1
OmniFlatten (Zhang et al., 2024a) 2025/1 2 -
RTTL-DG (Mai and Carson-Berndsen, 2025) 2025/1 2
Freeze-Omni Wang et al. (2024c) 2024/11 1 -
Closed-source Commercial Models
GPT-4o Voice Mode 2024/5 - - -
Gemini Live Voice Chat 2024/8 - - -
DouBao 2025/1 - - -
Table 1: Overview of full-duplex speech language models. “—” indicates properties that are not publicly available or unspecified. “E2E” denotes end-to-end speech modeling (without relying on text), “#ch” indicates the number of speech input channels, “Interrupt” refers to interruption handling, “BC” represents backchanneling capability, and “S2S Release” shows whether the complete speech-to-speech pipeline is publicly released (some models only release partial checkpoints without this functionality).
3.4 Data Curation
We explain how to curate the appropriate data for each evaluation dimension as follows:
  • Candor (Pause Handling, Smooth Turn-taking): We utilize Candor Reece et al. (2023), an 850-hour open-ended spontaneous speech conversation dataset with two-channel recording, as our data source. To select suitable data for our benchmark, for pause handling: we remove instances where the pause break within one turn exceeds 0.6 seconds but is shorter than 1.5 seconds. For smooth turn-taking, we calculate the gap between each pair of consecutive turns and retain only those where all gaps fall below 0.4 seconds. Since Candor does not offer an official train–test split, we randomly sample the dataset to 897 and 539 samples on pause handling and smooth turn-taking, respectively.
  • ICC (Backchannel): Umair et al. (2024) use the In Conversation Corpus (ICC) to collect Transition Relevance Places (TRPs), which are points in a speaker’s utterance that signal appropriate moments for the listener to respond. This dataset consists of 28.33 minutes of high-quality informal American English dialogues, capturing responses from 118 native speakers with no prior expertise in turn-taking research. Participants were instructed to provide brief backchannel responses (e.g., “hmm,” “yes”) whenever they found it appropriate. Their responses were recorded on individual audio channels, synchronized with the stimulus audio. On average, 59 participants responded to each stimulus turn, sometimes multiple times if they perceived multiple Transition Relevance Places, leading to an average of 159 responses per turn. This allows for estimating both the likelihood of perceiving a TRP at a given moment and the distribution of these perceived response locations. By segmenting the audio into 200-ms time windows and normalizing the backchannel counts, we generate a ground truth backchannel distribution $Q$ for computing JSD for each stimulus.
  • Synthetic Data (User Interruption, Pause Handling): To address the scarcity of user interruption in existing public datasets, we generated text dialogues containing both contextual and interrupting turns using GPT-4o (OpenAI, 2023). Prompts for generating these user interruption data are provided in Appendix Table 4. Speech synthesis was performed using ChatTTS [^2] https://github.com/2noise/ChatTTS is utilized to synthesize speech from text with 10 different speaker voices randomly assigned to each sample. Leveraging the pause break tag ([break]) supported by ChatTTS, we also generated synthetic pause handling data to evaluate the model’s robustness to pauses. The prompt for generating pause handling data is also found in Appendix Table 3. This process resulted in 200 samples for user interruption and 137 samples for pause handling.
Dimension Pause Handling Backchannel Smooth Turn Taking User Interruption
Data Synthetic Candor ICC Candor Synthetic
Metric TOR ($\downarrow$) TOR ($\downarrow$) TOR ($\downarrow$) Freq ($\uparrow$) JSD ($\downarrow$) TOR ($\uparrow$) Latency ($\downarrow$) TOR ($\uparrow$) GPT-4 ($\uparrow$) Latency ($\downarrow$)
dGSLM 0.949 0.953 0.782 0.013 0.950 0.989 0.572 0.895 0.201 3.972
Moshi 1.000 0.989 1.000 0.005 0.977 0.996 0.112 1.000 0.765 0.037
Freeze-Omni 0.672 0.287 0.782 0.002 0.984 0.369 1.168 0.795 3.371 1.200
Table 2: Models comparison. We evaluate several models across different conversational dimensions, where Latency is presented in seconds.
4 Models Under Evaluation
To fully assess full-duplex behavior, we review these models and compare their key features in Table 1. Some systems have detailed designs and are classified as Transparent Models (e.g., SyncLLM Veluri et al. (2024), Moshi Défossez et al. (2024), dGSLM Nguyen et al. (2023), VITA Fu et al. (2024), Freeze-Omni Wang et al. (2024c), SALMONN-omni Yu et al. (2024), Wang et al. (2024a), Parrot Wang et al. (2025), OmniFlatten Zhang et al. (2024a), MinMo Chen et al. (2025), and RTTL-DG Mai and Carson-Berndsen (2025)). In contrast, Closed-Source Commercial Models (e.g., GPT-4o Voice Mode [^3] https://openai.com/index/hello-gpt-4o/, Gemini Live Voice Chat [^4] https://blog.google/products/gemini/made-by-google-gemini-ai-updates/, and Doubao [^5] https://team.doubao.com/en/special/realtime_voice) are available only via apps or demos. For closed-source systems that claim full-duplex capabilities, we treat them as such despite limited design details.
Since many models do not release complete speech-to-speech checkpoints (see Table 1, column “S2S Release”), our evaluation focuses on open-source, full-duplex systems. By open-sourcing our data and metrics, we encourage the broader community, including those with closed-source models, to evaluate their systems using our benchmark. Currently, the models evaluated in this work are those with a publicly released speech-to-speech inference pipeline:
  • dGSLM Nguyen et al. (2023): dGSLM is a textless model designed to generate natural spoken dialogues directly from raw audio. Its three-stage architecture consists of: (1) a HuBERT+k-means encoder Hsu et al. (2021) that converts speech into discrete units; (2) a dual-tower Transformer-based dialogue language model with cross-attention for turn synchronization; and (3) a HiFi-GAN decoder that reconstructs speech waveforms. Trained on 2,000 hours of two-channel telephone conversations Cieri et al. (2004), dGSLM captures both verbal and non-verbal communication cues, such as laughter and backchanneling. While the original model lacks online interaction capabilities (as both two channels are predicted from context), we extend it to interact with live user speech, enabling to simulate the online user interaction. Details of this extension are provided in Appendix A.
  • Moshi  Défossez et al. (2024): Moshi treats spoken dialogue as a speech-to-speech generation task by combining a 7B text-based LLM (Helium) with a neural audio codec (Mimi) that uses residual vector quantization for speech processing. Its hierarchical design leverages an “Inner Monologue” to predict time-aligned text tokens before generating audio tokens, enhancing linguistic quality. Unlike dGSLM, Moshi processes real user speech in real time through a multi-stream design that supports overlaps and interruptions. While it serves as a general-purpose dialogue system, its complex inference process requires task-specific adaptations, which are detailed in Appendix B.
    Figure 3: Ablations on backchanneling. We vary the VAD Silero (2024) threshold for model response segmentation and evaluate backchanneling cues. The VAD threshold ranged from 0.1 to 0.6. For JSD, lower values indicate more natural responses, while lower TOR values are desirable. Higher frequency values indicate more backchannels. dGSLM consistently delivered more backchannel capabilities compared to other models.
  • Freeze-Omni (Wang et al., 2024c): Freeze-Omni is a cascaded full-duplex speech-to-speech dialogue model that achieves real-time interaction using a chunk-wise state prediction mechanism combined with a streaming speech architecture, while keeping the LLM frozen. The system utilizes a voice activity detector (VAD) to identify the start of speech and processes it in chunks using a streaming encoder. This encoder maps acoustic features into the LLM text space. A supplementary classification layer predicts the conversational state of each chunk, enabling real-time detection of whether the user is continuing the dialogue or initiating an interruption. To manage interruptions effectively, the system incorporates multiple modules functioning as monitors and speakers, facilitating full-duplex interaction. Local inference is performed using the official release server.
5 Results
Table 2 presents the results on four evaluation dimensions.
Avoiding Interruptions During Speaker Pauses
All three SDMs show a high Takeover Rate (TOR) when handling pauses, suggesting they often interrupt speakers during natural breaks. The end-to-end models (dGSLM and Moshi) tend to interrupt more frequently in both real and synthetic datasets. In contrast, Freeze-Omni, which includes a dedicated module for predicting speaking and listening states, produces a lower interruption rate. This indicates that a state-control component helps manage pauses more effectively, and it may benefit end-to-end systems if integrated.
Backchanneling Dynamics
We examined three models (dGSLM, Moshi, and Freeze-Omni) on the ICC corpus using TOR, backchannel frequency (Freq), and Jensen–Shannon Divergence (JSD). Similar to the pause-handling results, Moshi often takes over the turn with whisper-like responses, resulting in a high TOR. dGSLM and Freeze-Omni have lower TOR values, with dGSLM achieving the highest Freq and lowest JSD, indicating stronger backchannel behavior and a more natural conversational flow. Freeze-Omni remains largely silent, providing few backchannels.
We also tested different thresholds for the VAD Silero (2024), ranging from 0.1 to 0.6 (with a default of 0.5), to see how speech-segmentation granularity affects model performance (Table 2 and Figure 3). dGSLM shows the greatest variation, especially in Freq and JSD, while Moshi continues to speak through most turns (TOR = 1.0). Freeze-Omni is relatively stable but produces fewer backchannels. This confirms that despite segment granularity affects model performance, each model’s core tendencies remain: dGSLM backchannels actively, Moshi whispers persistently, and Freeze-Omni stays mostly silent.
Latency of Turn-Taking
We measured TOR and average response times on the Candor dataset. Both dGSLM and Moshi produce high TOR values and respond quickly. Freeze-Omni shows a higher latency due to its cascaded design, which first generates text and then uses text-to-speech. Its lower TOR may reflect missed opportunities to take over the turn because it fails to detect the end of a speaker’s segment.
Managing User Interruptions
Freeze-Omni achieves the best performance in user interruption handling, with significantly higher contextual relevance scores and acceptable latency. This is likely due to its robust large language model and multi-module architecture, which help manage user-initiated interruptions. By contrast, the end-to-end SDMs produce less coherent responses. Moshi responds quickly but yields a lower GPT-4 score, while dGSLM scores poorly overall, with high latency and reduced content quality. These results suggest that fully end-to-end systems face additional challenges in maintaining semantic coherence when users interrupt.
6 Conclusion
In this paper, we introduced Full-Duplex-Bench, a new benchmark designed to evaluate key aspects of full-duplex spoken dialogue models. By focusing on critical dimensions (pause handling, backchanneling, smooth turn-taking, and user interruption management), our framework fills an important gap left by existing benchmarks that largely target half-duplex systems or rely on coarse corpus-level metrics.
Our evaluation protocol employs automatic metrics to enable systematic and reproducible assessments of real-time interaction. Experiments on open-source full-duplex models reveal both promising capabilities and clear areas for improvement, particularly in achieving natural turn transitions and appropriate response timing.
We believe that Full-Duplex-Bench will serve as a valuable resource for the research community. By open-sourcing our data and evaluation metrics, we hope to foster collaboration and drive the development of more interactive and human-like spoken dialogue systems.
Limitations and Future Works
While our benchmark marks an important step in evaluating full-duplex SLMs, it also reveals several limitations that we plan to address in future work:
  • English-Only Corpora: Currently, our benchmark data is in English. However, turn-taking behavior may vary across languages and cultures (Stivers et al., 2009), and backchanneling signal words and frequency might differ between languages (White, 1989).
  • Closed-Sourced Commercial Model: We did not evaluate the commercial models, as it only allows interaction via a microphone and does not support local inference. Reproducing results is challenging due to variations in recording environments. In the future, we plan to explore better ways to interact with such models while ensuring reproducibility and efficiency.
  • More Accurate Backchannel Detection: We currently detect backchannels based on short duration and minimal word usage. However, incorporating a more advanced detection model that integrates temporal, lexical, and prosodic signals could improve accuracy.
  • Non-lexical Expression: Spontaneous spoken dialogue includes non-lexical expressions such as laughter and sighs. While our approach transcribes speech into text using ASR, our current metrics and evaluation dimensions do not account for these non-lexical elements. Future work should incorporate such expressions to enhance the interactivity and naturalness of full-duplex models.
Appendix A dGSLM Implementation
Denote user speech $U=[U_{1},U_{2},...,U_{T}]$ as the user stream and $M=[M_{1},M_{2},...,M_{T}]$ as the model stream, where $U_{i}$, $M_{i}$ are speech tokens after tokenization (Hsu et al., 2021), and $i=1,2,...,T$. In vanilla dGSLM (Nguyen et al., 2023), the model takes a two-channel prompt $([U_{1},U_{2},...,U_{p}],[M_{1},M_{2},...,M_{p}])$ and samples full sequences $([U_{1},U_{2},...,U_{p},U_{p+1},...,U_{T}],$ $[M_{1},M_{2},...,M_{p},M_{p+1},...,M_{T}])$ in an cross modality autoregressive manner. Specifically, $U_{i},M_{i}=f_{\theta}(U_{<i},M_{<i})$. Note that under this setting, it is not straightforward to achieve online human-model interaction, as user speech $U$ is generated and then fed as input. Based on this context, we extend dGSLM to allow flexible speech interaction, and we detail each variant below. The overall paradigm is visualized in Figure 4.
dGSLM Pause Handling/Backchanneling
We apply user speech as teacher forcing and only generate model speech. That is, $M_{i}=f_{\theta}(U^{T}_{<i},M_{<i})$, where $U^{T}$ denotes ground truth user speech. Different from vanilla dGSLM, all user speech serves as input and all model speech is output. We then analyze model output speech, which is still time synchronous with user speech, to evaluate pause handling and backchanneling performance. We also detail our teacher forcing algorithm in Sec. 21. $S_{U,t},S_{M,t}$ represent scores for user and model streams at time $t$. [BOS], [EOS] represent beginning and end of sequence tokens
Algorithm 1 Teacher Forcing Generation with dGSLM
1:User speech tokens $U^{T}$, beam size $k$, max length $T$
2:Generated model speech tokens $M$
3:Initialize $M$ with padding tokens
4:Initialize scores matrix $S$ with zeros
[…]
6:for $t=1$ to $T$ do
7:     $\textit{lprobs}_{U},\textit{lprobs}_{M}\leftarrow f_{\theta}(U^{T}_{<t},M_{<t})$triangleright Get logits for both streams
8:     if $t<|U^{T}|$ then
9:         $U_{t}\leftarrow U^{T}_{t}$triangleright Teacher forcing user stream
10:         $S_{U,t}\leftarrow 0$triangleright Zero score for teacher forced tokens
11:     else
[…]
13:     end if
14:     $\textit{cand}_{M}\leftarrow\text{BeamSearch}(\textit{lprobs}_{M},k)$triangleright Search for model tokens
15:     $M_{t}\leftarrow\text{TopBeam}(\textit{cand}_{M})$triangleright Select top beam candidate
16:     $S_{M,t}\leftarrow\text{Score}(\textit{cand}_{M})$triangleright Update model scores
[…]
18:         break
19:     end if
20:end for
21:return $(M,S)$triangleright Return generated sequence and scores
dGSLM Smooth Turn Taking
Evaluation of turn-taking is implemented as a simple spoken QA task where we process user speech and await model responses. To integrate such spoken QA capability in dGSLM, we first define prefix tokens $([U_{1},U_{2},...,U_{p}],[M_{1},M_{2},...,M_{p}])$, where the user branch contains user speech while the model branch contains pure silence. We then process a new user prompt as input $[U^{T}_{p+1},...,U^{T}_{T}]$ consisting of pure silence, and apply the same teacher forcing algorithm (Sec. 21) to generate the model response: $M_{i}=f_{\theta}(U_{1:p},M_{1:p},U_{<i}^{T})$.
dGSLM User Interruption
We implement user interruption using two dGSLM modules. As shown in Figure 4, in the first stage, we apply vanilla dGSLM inference where, given a user context wav and a silent model wav, we prompt both user and model responses, keeping only the model response while setting the user response to silence. When integrating user-interrupting speech into the current conversational stream, we first adopt offline VAD Silero (2024) to detect model responses. We select the first significant segment (duration > 1s) and obtain its boundary $[T_{\text{start}},T_{\text{end}}]$. We inject the interrupting voice at time $T_{I}=(T_{\text{end}}-T_{\text{start}})/2$. Then we concatenate all historical information and form a two-channel speech prompt. Let the user context token length be $p$ and the user interrupting token length be $q$. Our user stream will be $[U_{1},...,U_{p}]$, $[U_{p+1},...,U_{T_{m}}]$, $[U_{T_{m}},...,U_{T_{m}+q},...,U_{T}]$, corresponding to context speech, padded silence, user interruption speech, and padded silence if necessary. Our model stream will be $[M_{1},...,M_{T_{\text{start}}}]$, $[M_{T_{\text{start}}+1},...,M_{T_{\text{end}}}]$, $[M_{T_{\text{end}}+1},...,M_{T}]$, corresponding to silent speech, model’s first significant response, and remaining speech (if there is more than one significant response or backchannels), with padded silence if necessary. The second dGSLM module takes the concatenated prompts $[U_{1},...,U_{T}]$ and $[M_{1},...,M_{T}]$ as input and generates a subsequent response sequence $[M_{T+1},...,M_{T^{\prime}}]$, where we retain only the model’s generated response for evaluation purposes
Figure 5: Moshi’s inference pipeline for different conversational scenarios. (a) Vanilla time-asynchronous response generation with prefilling and silent response generation, which is also the pipeline that the smooth turn taking follows. (b) Time-synchronous generation for backchanneling and pause handling, and (c) User interruption handling through controlled user speech timing, where silence and user speech are strategically input to generate appropriate model responses. All scenarios utilize a single model with 80ms chunk-wise processing.
Appendix B Moshi Implementation
Given user speech $U=[U_{1},U_{2},...,U_{T}]$, Moshi Défossez et al. (2024) predicts the model stream $M=[M_{1},M_{2},...,M_{T}]$ in a streaming manner (chunk-wise, with each chunk spanning 80ms), enabling conversational dynamics such as backchannels. Like dGSLM, Moshi can also predict the user stream, providing flexibility for general-purpose spoken dialogue systems. We detail Moshi’s adaptation to different inference tasks in Figure 5.
Vanilla Design
The fundamental task in spoken dialogue is time-asynchronous speech-to-speech response. To achieve this, we input user speech $U$ into the system. Although Moshi will also autoregressively generate time-synchronous responses in real-time, this part is ignored during Prefilling. After prefilling, silence speech is used as input, and Moshi autoregressively generates its response. In this case, there is no speech overlap. Smooth Turn Taking follows this vanilla pipeline.
Moshi Backchanneling and Pause Handling
Time-synchronous model response generation is required to explore backchanneling and pause-handling capabilities. This process is straightforward: we input speech, and the model autoregressively generates the synchronous response.
Moshi User Interrupting
User interruption modeling requires fine-grained control of user speech timing. To achieve this, we first perform prefilling (inputting user speech while ignoring model output). Unlike vanilla or smooth turn-taking designs, instead of inputting the entire silence speech user stream to generate a response, we predetermine a generation length. For the first half of the generation, we input silence; for the second half, we begin inputting the entire user interrupt speech until reaching the maximum generation length. If any user interrupts speech remains unprocessed, we feed it into the model to continue generating model speech. Finally, after fully processing the user’s interrupting speech, we input silence as a prompt to generate the model response. In contrast to dGSLM’s design (Sec. A), this process uses only one model.
Appendix C Additional Backchanneling Ablation
We also reported the standard deviation of each metric across three models. The results are displayed in Table 6. We also visualize these statistics in Figure 6.
Figure 6: VAD ablation study. We vary the VAD Silero (2024) threshold for the backchanneling behaviors of several models.
Model TOR Freq JSD
dGSLM 0.782 ± 0.413 0.013 ± 0.028 0.950 ± 0.028
Moshi 1.000 ± 0.000 0.005 ± 0.015 0.977 ± 0.067
Freeze-Omni 0.782 ± 0.413 0.002 ± 0.007 0.984 ± 0.058
Table 6: Performance comparison of models on backchanneling metrics